OpenStudy (anonymous):

Correction to my original problem: 36x^3 + x^2 - 21x = 0 I had it set to 2nd degree at first. Needs to be factored to solve.

OpenStudy (radar):

x(36x^2 + x - 21) x(9x+7)(4x-3)=0 x=0 x=-7/9 x=3/4

OpenStudy (anonymous):

How did you get the second line? I see that they are broken down from 36x^2 and - 21, but I thought that you have to solve by getting the product of 36 and 21 and find number that multiply to the product and add to 1. Thanks

OpenStudy (radar):

Your thoughts were correct the product of 9 and 4 is 36. The combination of (28x and -27x) does = x . the 28x is the result of 7 times 4x, the -27x is the result of 9x times -3.

OpenStudy (anonymous):

Yeah, I figured it out right after I asked you. Thank you for your help though. I really appreciate it!